How fast would a vehicle going 260 mph reach an 80 mile checkpoint
Answer by stanbon(75887)

How fast would a vehicle going 260 mph reach an 80 mile checkpoint?
--------
time = distance/rate
t = 80/260 = 4/13 hour = 10 min 27.6 sec
--------------
Cheers,
Stan H.
----------------